OpenStudy (anonymous):

if sin30 DEGREES=1/2, FIND CSC30DEGREES

OpenStudy (anonymous):

if \(\sin(30)=\frac{1}{2}\)??

sam (.sam.):

\[\csc(30)\] \[=\frac{1}{\sin(30)}\] \[=\frac{1}{\frac{1}{2}}\] =2

sam (.sam.):

Just use \[\csc(x)=\frac{1}{\sin(x)}\]